Solve 4 sin ) = 3 for the four smallest positive solutic
Answer #1


4 sin ( x ) = 3 - 2 » Sin x) = 2 = 0.75 = 48. sulu8.580) Sin & n) = sin (0.26982) 186 = radian to - A radian 48.590 K x 48.59